### v2

@@ 1 Introduction
Finding correspondences between deformable shapes is a fundamental problem in geometry processing, with various applications such as texture and deformation transfer [dinh2005texture, ovsjanikov2012functional], or statistical shape analysis [hasler2009statistical, bogo2014faust, elamrani2024fuss]. The problem has been widely studied in the literature  [ovsjanikov2012functional, eisenberger2020, abdelreheem2023zero, wimmer2024back, cao2023unsupervised], and can vary significantly in complexity. For example, current methods still struggle when estimating severe non-isometric deformations for shapes of different categories [dyke2020shrec, panine2022non, eisenberger2020, bastian2023hybrid, cao2024spectral], or when they are only partially observed [roetzer2022scalable, ehm2024partial]. Current non-isometric shape datasets are often limited in size and flexibility [Zuffi:CVPR:2017, melzi2019shrec, dyke2020shrec, magnet2022smooth, li20214dcomplete], restricting their ability to support dense correspondence propagation or custom annotations, which reduces their applicability to broader tasks.
In real-world scenarios, it is particularly common to encounter shapes with missing parts due to occlusions, limited sensor coverage, or acquisition errors during 3D scanning. The setting in which at least part of one shape does not exist on the other is called partial shape matching [rodola2017partial, attaiki2021dpfm, roetzer2022scalable, ehm2023geometrically, ehm2024partial]. Despite its practical relevance, partial shape matching has received limited attention in the context of machine learning-based approaches, partly because of the increased difficulty, but largely due to the lack of suitable and realistic large-scale datasets, a requirement for data-hungry learning algorithms. The full-to-full matching setting is not only simpler but also benefits from wider data availability which has supported the development of data-driven methods that already display convincing performance, even without supervision [halimi2019unsupervised, roufosse2019unsupervised, cao2023unsupervised]. In contrast, existing datasets for partial shape correspondence often suffer from a small number of shapes, limited realism of generated partiality (e.g., cuts or holes) [cosmo2016shrec], and a lack of diversity in shape categories [bracha2023partial, attaiki2021dpfm]. This limited data availability hampers the progress of data-driven methods for partial shape matching; recent approaches have thus required pre-training with large full-geometry datasets [cao2023unsupervised], or rely on ground-truth correspondences [attaiki2021dpfm] which are challenging to obtain in practice.
@@ 2.1 Datasets
Full-to-Full Datasets. Full-to-full shape correspondence datasets, such as FAUST [bogo2014faust] and SCAPE [anguelov2005scape], have been essential in methodological progress on shape correspondence. They encompass humans in various poses, initially emphasising point-wise correspondence recovery under isometric deformations. SHREC’19 [melzi2019shrec] was introduced to evaluate the generalisation capabilities of methods to different mesh connectivity on human scans. Datasets like TOSCA [bronstein2008numerical], SMAL [Zuffi:CVPR:2017], SHREC’20 [dyke2020shrec], and DeformingThings4D [magnet2022smooth, li20214dcomplete] have shifted the emphasis towards non-isometric correspondence between various animal categories and humanoid figurines. Although deep learning methods have achieved state-of-the-art performance in shape correspondence, non-isometric datasets remain relatively small and static, limiting their applicability in training and evaluating deep learning-based correspondence methods. While full-to-full shape matching has received considerable attention, the partial setting has been explored less (see Table 1).
@@
[…]
Partial-to-Full Datasets. In this setting, a partial shape is deformably matched to a shape template with complete geometry. SHREC’16 [cosmo2016shrec], based on the TOSCA [bronstein2008numerical] dataset, was the first benchmark for this problem setting. PFAUST [bracha2023partial] and PFARM [attaiki2021dpfm] were later introduced using a similar approach based on the FAUST [bogo2014faust] and FARM [kirgo2021wavelet] datasets, respectively. SHREC’20 [dyke2020shrec] also contains a subset of full-to-partial correspondence pairs.
Partial-to-Partial Datasets. Partial-to-partial shape matching involves matching two partial shapes with unknown overlap, making it considerably more challenging than partial-to-full matching. In [attaiki2021dpfm] the authors proposed the first benchmark for this setting with CP2P [attaiki2021dpfm] based on the SHREC’16 CUTS [cosmo2016shrec] benchmark. Recently, PSMAL [ehm2024partial] was introduced, extending the SMAL [Zuffi:CVPR:2017] dataset to include partial-to-partial correspondence pairs.
@@ 2.2 Methods for Shape Correspondence
We proceed by discussing a selection of shape matching methods and refer to recent surveys [sahillioglu_recent_2020, deng_survey_2022] for a more thorough treatment. While point-cloud matching methods also exist [li2022lepard, bai2020d3feat, huang2021predator, yu2023rotation], we focus on mesh-based methods, which are better aligned with the surface representations emphasised in our dataset. Methods addressing shape correspondence can be categorised into axiomatic and learning-based. Among the axiomatic approaches are discrete optimization methods [bernard2020mina] and spectral methods utilising the powerful framework of functional maps [ovsjanikov2012functional, rodola2017partial, melzi2019zoomout]. However, the performance of the latter class typically degrades as assumptions of isometry are violated, such as in partial correspondence settings.
@@
Spectral methods based on the eigenfunctions of the Laplace-Beltrami Operator (LBO) have been successful in full-to-full matching settings [ovsjanikov2012functional, litany2017deep, donati2020deepGeoMaps, bastian2023hybrid, cao2023unsupervised], yielding a compact solution. In contrast to spectral approaches, discrete optimization formulations have also been introduced for full-to-full matching [bernard2020mina, roetzer2022scalable, roetzer2023fast, roetzer2024spidermatch]. These algorithms often come with guaranteed bounds on optimality, sometimes incorporating geometric consistency constraints.
@@
Rodola et al. [rodola2017partial] demonstrate that for partial-to-full matching, the functional map has a slanted diagonal structure, which has, for example, been used in [litany2016non] to solve multi-part partial-to-partial correspondence; however, under the assumption that a full template is given. Several methods also operate in the spectral domain in learning-based settings [attaiki2021dpfm, cao2022unsupervised, cao2023unsupervised]. Other works do not use functional map layer [bracha2023partial, bracha2024wormhole] and explicitly tackle the partial-to-full case.
@@
Few works thus far explicitly address the partial-to-partial matching setting, which is particularly challenging as the overlapping region between shapes is unknown. As such, the overlapping region must simultaneously be identified and matched under an unknown deformation. DPFM [attaiki2021dpfm] demonstrates that spectral methods are applicable in this domain using an additional attention mechanism, albeit using ground-truth supervision for feature learning. In [windheuser2011geometrically] shape matching is formulated as an integer linear program (ILP) over the space of orientation-preserving diffeomorphisms. SM-COMB [roetzer2022scalable] demonstrated that the latter constitutes a strong prior to guide partial-to-partial matching. However, holes must be synthetically filled to obtain a closed manifold. Recently, GC-PPSM [ehm2024partial] formulated a geometrically consistent partial-to-partial shape matching non-linear integer program. While both SM-COMB and GC-PPSM are compatible with various surface descriptors, they achieve SOTA results only with high-quality pre-trained input features (e.g. from ULRSSM [cao2023unsupervised]), which can be prohibitive when these are not available [roetzer2022scalable, ehm2024partial].
@@ 2.3 Feature Descriptors for Shape Correspondence
Many axiomatic and learning-based methods for shape matching rely on surface feature descriptors as guidance for correspondence computation. Feature descriptors, like SHOT [tombari2010unique], can be handcrafted based on extrinsic quantities like XYZ-coordinates or surface normals. Many surface descriptors are sensitive to surface degradation, such as holes, or are not invariant under rotations, as is the case for XYZ coordinates. The common intrinsic heat kernel (HKS) [sun_concise_2009] and wave kernel signatures (WKS) [aubry_wave_2011] are invariant under isometric deformations; however, they are not effective under partiality as the isometry assumption can be heavily violated. On the other hand, methods like DiffusionNet [sharp2022diffusionnet] aim to learn robust surface descriptors from data while still using some features like XYZ coordinates or WKS descriptors as input. Considering that 3D shapes of many datasets are typically oriented in an upright manner [bogo2014faust, anguelov2005scape, bronstein2008numerical, rodola2014dense, dyke2020shrec], or generally positioned in a canonical pose [eisenberger2021neuromorph] this can lead to biased predictions. We discuss these biases in detail in Sec. 5.4. Moreover, learned features require large training datasets and a high-quality (self-)supervision signals which is challenging to achieve under severe partiality. Since obtaining per-vertex feature descriptors directly on 3D shapes is challenging, [wei2016dense] addresses this by establishing correspondences between 3D shapes through correspondences between their 2D rendered images. Recent methods [dutt2024diff3f, decatur20233d, kobayashi2022decomposing, fischer2024nerf] leverage 2D image foundation model features [oquab2023dinov2, rombach2022high] from multiple rendered views of a 3D shape to aggregate per-vertex features. However, such an approach can suffer from significant bias due to missing semantic information in the rendered views.
@@ 3.1 Cross-Dataset/Category Correspondences
Current datasets commonly lack correspondences between different categories, limiting data availability and diversity [bronstein2008numerical, li20214dcomplete]. We address this shortcoming by manually establishing correspondences across categories and across different datasets. To this end, we treat each category for which a cross-category correspondence is not available due to missing annotation as a separate shape group. Then, we select a template shape for each group and manually choose a set of corresponding landmarks on pairs of templates using FaceForm [faceform] (see supp. mat. Sec. 9). In the case of non-isometry, more manual landmarks are needed to guide the deformation, see Fig. 10 in supp. mat. for an example. We then morph one template into another to establish dense point-to-surface correspondences (represented in barycentric coordinates to account for different discretisations). Using the tightly aligned shapes, dense correspondences are obtained by projecting points onto triangles of the other surface. After the templates are in dense correspondence, we propagate them to the entire shape category using the existing intra-category correspondences. Overall, we build a shape network of pairwise dense correspondences that cover all possible shape templates (cross-category and cross-dataset). The detailed shape network is presented on our project page. A colourmap visualization of the correspondences can be seen in Fig. 1.
@@ 3.3 Alignment of Shapes
Many correspondence datasets, specifically partial ones, such as CP2P [attaiki2021dpfm] and SHREC’16 [cosmo2016shrec], keep shapes in the same canonical pose during training and testing. This bias can prevent methods from generalising to shapes under random rigid transformations, see also [attaiki2023generalizable]. As this assumption is often unrealistic in real-world settings, we randomly rotate the shapes around the z-axis (with respect to their canonical pose as present in the original dataset) and thus generate shapes in random upright positions. We show an ablation study about rotation in Sec. 5.4.
@@ 3.4 Custom Annotation Propagation
Shape annotations like movable parts labels used in robotics [qiu2025articulate] or part segmentation labels [kalogerakis2010learn, chen2024DAE-Net] are often indispensable for data-driven training or evaluation. Our framework allows for the propagation of any type of information defined on the vertices from a few annotated examples to all shapes in the dataset. By leveraging the shape network, we can find a path to the closest annotated example from each shape, establishing correspondences to a shape that offers ground-truth information. To demonstrate the effectiveness of our method, we choose to propagate left/right annotations from only $12$ manually annotated shapes to the whole dataset (see Figure 2). While in theory, even fewer annotations would suffice, we selected $12$ shapes to reduce propagation error (see the examples in the supplementary material in Figure 8).
@@ 3.5 Partiality Generation
Flexible Generation of Shape Matching Instances. Existing partial shape datasets provide a precomputed and fixed number of holes/cuts [cosmo2016shrec, bracha2023partial, kirgo2021wavelet, attaiki2021dpfm, dyke2020shrec, ehm2024partial], creating partial shape pairs from identical meshes in varying poses. Instead, our proposed framework can generate partial shapes from a larger and more diverse pool of shapes, allowing for pairing compatible shapes from different categories, potentially originating from different datasets. This creates a more challenging scenario encompassing simultaneous non-isometric deformations and partiality, representing an important step towards more realistic settings. Furthermore, our framework introduces more variety by using different camera poses for ray casting and re-meshing shapes to different resolutions.
@@ 3.6 Semantic Missing Parts
By combining multiple categories and datasets, we encounter cases where dense correspondences cannot be reliably established between certain parts of shapes without introducing inaccuracies (e.g., matching a deer’s horn to a horse’s head). While such ambiguities can be managed when considering semantic correspondences [abdelreheem2023zero], they remain problematic for dense correspondences. To address this, for the first time we create large cross-dataset correspondences with semantic missing parts annotations, as shown in Fig. 3. To this end, we modify our pipeline described in Sec. 3.1 by manually removing the semantic missing parts from the shapes and match the rest of the shapes to get the dense correspondences. We show an example in Fig. 10 in supp. mat. This creates a novel objective for shape matching methods that warrants future research.

### v3

@@ 1 Introduction
Finding correspondences between deformable shapes is a fundamental problem in geometry processing, with various applications such as texture and deformation transfer [DYT05, OBS*12], or statistical shape analysis [HSS*09, BRLB14, ECB24]. The problem has been widely studied in the literature  [OBS*12, ELC20, AEOW23, WWO24, CRB23], and can vary significantly in complexity. For example, current methods still struggle when estimating severe non-isometric deformations for shapes of different categories [DLR*20, PKO22, ELC20, BXNL24, CEE*24], or when they are only partially observed [RSCB22, EGR*24]. Current non-isometric shape datasets are often limited in size and flexibility [ZKJB17, MMR*19, DLR*20, MRSO22, LTT*21], restricting their ability to support dense correspondence propagation or custom annotations, which reduces their applicability to broader tasks.
In real-world scenarios, it is particularly common to encounter shapes with missing parts due to occlusions, limited sensor coverage, or acquisition errors during 3D scanning. The setting in which at least part of one shape does not exist on the other is called partial shape matching [RCB*17, APO21, RSCB22, ERE*23, EGR*24]. Despite its practical relevance, partial shape matching has received limited attention in the context of machine learning-based approaches, partly because of the increased difficulty, but largely due to the lack of suitable and realistic large-scale datasets, a requirement for data-hungry learning algorithms. The full-to-full matching setting is not only simpler but also benefits from wider data availability which has supported the development of data-driven methods that already display convincing performance, even without supervision [HLR*19, RSO19, CRB23]. In contrast, existing datasets for partial shape correspondence often suffer from a small number of shapes, limited realism of generated partiality (e.g., cuts or holes) [CRB*16], and a lack of diversity in shape categories [BDK23, APO21]. This limited data availability hampers the progress of data-driven methods for partial shape matching; recent approaches have thus required pre-training with large full-geometry datasets [CRB23], or rely on ground-truth correspondences [APO21] which are challenging to obtain in practice.
@@ 2.1 Datasets
Full-to-Full Datasets. Full-to-full shape correspondence datasets, such as FAUST [BRLB14] and SCAPE [ASK*05], have been essential in methodological progress on shape correspondence. They encompass humans in various poses, initially emphasising point-wise correspondence recovery under isometric deformations. SHREC’19 [MMR*19] was introduced to evaluate the generalisation capabilities of methods to different mesh connectivity on human scans. Datasets like TOSCA [BBK08], SMAL [ZKJB17], SHREC’20 [DLR*20], and DeformingThings4D [MRSO22, LTT*21] have shifted the emphasis towards non-isometric correspondence between various animal categories and humanoid figurines. Although deep learning methods have achieved state-of-the-art performance in shape correspondence, non-isometric datasets remain relatively small and static, limiting their applicability in training and evaluating deep learning-based correspondence methods. While full-to-full shape matching has received considerable attention, the partial setting has been explored less (see Table 1).
@@
F2F FAUST [BRLB14] 100 [Uncaptioned image]
SCAPE [ASK*05] 71
SMAL [ZKJB17] 49 [Uncaptioned image]
DT4D Human [MRSO22, LTT*21] 248
DT4D Animal [MRSO22, LTT*21] 1950
TOSCA Human [BBK08] 45
TOSCA Four-legged [BBK08] 37
KIDS [RRW*14] 32
SHREC’20 [DLR*20] 11
SHREC’19 [MMR*19] 44
P2F SHREC’16 [BBK08, CRB*16] 76 /
PFAUST [BDK23, BRLB14] 10
FARM [KMP*21] 5
PFARM [APO21, KMP*21] 28
SHREC’20 [DLR*20] 14
P2P CP2P [APO21, CRB*16] 76 /
PSMAL [EGR*24, ZKJB17] 43
Both BeCoS (ours) 2543 /
Table 1: An Overview of 3D Deformable Shape Datasets, categorised as full-to-full (F2F), partial-to-full (P2F) and partial-to-partial (P2P), including the number of available shapes, shape type (Humanoid or Four-legged ), and deformation class (Near-Isometric () and Non-Isometric ()). Due to limited correspondence pairs, FARM and PFARM are typically only used for evaluation. Note that in previous datasets, not all unique shapes have annotated correspondences, limiting the number of possible correspondence pairs, e.g., DT4D Animal (maximum 62 452 correspondence pairs) compared to our benchmark (maximum 1 414 179 correspondence pairs). As SHREC20 [DLR*20] contains three partial shapes, it can be used as a partial-to-full or full-to-full dataset.
Partial-to-Full Datasets. In this setting, a partial shape is deformably matched to a shape template with complete geometry. SHREC’16 [CRB*16], based on the TOSCA [BBK08] dataset, was the first benchmark for this problem setting. PFAUST [BDK23] and PFARM [APO21] were later introduced using a similar approach based on the FAUST [BRLB14] and FARM [KMP*21] datasets, respectively. SHREC’20 [DLR*20] also contains a subset of full-to-partial correspondence pairs.
Partial-to-Partial Datasets. Partial-to-partial shape matching involves matching two partial shapes with unknown overlap, making it considerably more challenging than partial-to-full matching. In [APO21] the authors proposed the first benchmark for this setting with CP2P [APO21] based on the SHREC’16 CUTS [CRB*16] benchmark. Recently, PSMAL [EGR*24] was introduced, extending the SMAL [ZKJB17] dataset to include partial-to-partial correspondence pairs.
@@ 2.2 Methods for Shape Correspondence
We proceed by discussing a selection of shape matching methods and refer to recent surveys [Sah20, DYDZ22] for a more thorough treatment. While point-cloud matching methods also exist [LH22, BLZ*20, HGU*21, YQH*23], we focus on mesh-based methods, which are better aligned with the surface representations emphasised in our dataset. Methods addressing shape correspondence can be categorised into axiomatic and learning-based. Among the axiomatic approaches are discrete optimization methods [BST20] and spectral methods utilising the powerful framework of functional maps [OBS*12, RCB*17, MRR*19]. However, the performance of the latter class typically degrades as assumptions of isometry are violated, such as in partial correspondence settings.
@@
Spectral methods based on the eigenfunctions of the Laplace-Beltrami Operator (LBO) have been successful in full-to-full matching settings [OBS*12, LRR*17, DSO20, BXNL24, CRB23], yielding a compact solution. In contrast to spectral approaches, discrete optimization formulations have also been introduced for full-to-full matching [BST20, RSCB22, RAC*24, RB24]. These algorithms often come with guaranteed bounds on optimality, sometimes incorporating geometric consistency constraints.
@@
Rodola et al. [RCB*17] demonstrate that for partial-to-full matching, the functional map has a slanted diagonal structure, which has, for example, been used in [LRB*16] to solve multi-part partial-to-partial correspondence; however, under the assumption that a full template is given. Several methods also operate in the spectral domain in learning-based settings [APO21, CB22, CRB23]. Other works do not use functional map layer [BDK23, BDK24] and explicitly tackle the partial-to-full case.
@@
Few works thus far explicitly address the partial-to-partial matching setting, which is particularly challenging as the overlapping region between shapes is unknown. As such, the overlapping region must simultaneously be identified and matched under an unknown deformation. DPFM [APO21] demonstrates that spectral methods are applicable in this domain using an additional attention mechanism, albeit using ground-truth supervision for feature learning. In [WSSC11] shape matching is formulated as an integer linear program (ILP) over the space of orientation-preserving diffeomorphisms. SM-COMB [RSCB22] demonstrated that the latter constitutes a strong prior to guide partial-to-partial matching. However, holes must be synthetically filled to obtain a closed manifold. Recently, GC-PPSM [EGR*24] formulated a geometrically consistent partial-to-partial shape matching non-linear integer program. While both SM-COMB and GC-PPSM are compatible with various surface descriptors, they achieve SOTA results only with high-quality pre-trained input features (e.g. from ULRSSM [CRB23]), which can be prohibitive when these are not available [RSCB22, EGR*24].
@@ 2.3 Feature Descriptors for Shape Correspondence
Many axiomatic and learning-based methods for shape matching rely on surface feature descriptors as guidance for correspondence computation. Feature descriptors, like SHOT [TSD10], can be handcrafted based on extrinsic quantities like XYZ-coordinates or surface normals. Many surface descriptors are sensitive to surface degradation, such as holes, or are not invariant under rotations, as is the case for XYZ coordinates. The common intrinsic heat kernel (HKS) [SOG09] and wave kernel signatures (WKS) [ASC11] are invariant under isometric deformations; however, they are not effective under partiality as the isometry assumption can be heavily violated. On the other hand, methods like DiffusionNet [SACO22] aim to learn robust surface descriptors from data while still using some features like XYZ coordinates or WKS descriptors as input. Considering that 3D shapes of many datasets are typically oriented in an upright manner [BRLB14, ASK*05, BBK08, RRW*14, DLR*20], or generally positioned in a canonical pose [ENK*21] this can lead to biased predictions. We discuss these biases in detail in Sec. 5.4. Moreover, learned features require large training datasets and a high-quality (self-)supervision signals which is challenging to achieve under severe partiality. Since obtaining per-vertex feature descriptors directly on 3D shapes is challenging, [WHC*16] addresses this by establishing correspondences between 3D shapes through correspondences between their 2D rendered images. Recent methods [DMM24, DLH23, KMS22, FLN*24] leverage 2D image foundation model features [ODM*23, RBL*22] from multiple rendered views of a 3D shape to aggregate per-vertex features. However, such an approach can suffer from significant bias due to missing semantic information in the rendered views.
@@ 3.1 Cross-Dataset/Category Correspondences
Current datasets commonly lack correspondences between different categories, limiting data availability and diversity [BBK08, LTT*21]. We address this shortcoming by manually establishing correspondences across categories and across different datasets. To this end, we treat each category for which a cross-category correspondence is not available due to missing annotation as a separate shape group. Then, we select a template shape for each group and manually choose a set of corresponding landmarks on pairs of templates using FaceForm [Fac23] (see supp. mat. Sec. 9). In the case of non-isometry, more manual landmarks are needed to guide the deformation, see Fig. 10 in supp. mat. for an example. We then morph one template into another to establish dense point-to-surface correspondences (represented in barycentric coordinates to account for different discretisations). Using the tightly aligned shapes, dense correspondences are obtained by projecting points onto triangles of the other surface. After the templates are in dense correspondence, we propagate them to the entire shape category using the existing intra-category correspondences. Overall, we build a shape network of pairwise dense correspondences that cover all possible shape templates (cross-category and cross-dataset). The detailed shape network is presented on our project page. A colourmap visualization of the correspondences can be seen in Fig. 1.
@@ 3.3 Alignment of Shapes
Many correspondence datasets, specifically partial ones, such as CP2P [APO21] and SHREC’16 [CRB*16], keep shapes in the same canonical pose during training and testing. This bias can prevent methods from generalising to shapes under random rigid transformations, see also [ALO23]. As this assumption is often unrealistic in real-world settings, we randomly rotate the shapes around the z-axis (with respect to their canonical pose as present in the original dataset) and thus generate shapes in random upright positions. We show an ablation study about rotation in Sec. 5.4.
@@ 3.4 Custom Annotation Propagation
Shape annotations like movable parts labels used in robotics [QYW*25] or part segmentation labels [KHS10, CCZZ24] are often indispensable for data-driven training or evaluation. Our framework allows for the propagation of any type of information defined on the vertices from a few annotated examples to all shapes in the dataset. By leveraging the shape network, we can find a path to the closest annotated example from each shape, establishing correspondences to a shape that offers ground-truth information. To demonstrate the effectiveness of our method, we choose to propagate left/right annotations from only $12$ manually annotated shapes to the whole dataset (see Figure 2). While in theory, even fewer annotations would suffice, we selected $12$ shapes to reduce propagation error (see the examples in the supplementary material in Figure 8).
@@ 3.5 Partiality Generation
Flexible Generation of Shape Matching Instances. Existing partial shape datasets provide a precomputed and fixed number of holes/cuts [CRB*16, BDK23, KMP*21, APO21, DLR*20, EGR*24], creating partial shape pairs from identical meshes in varying poses. Instead, our proposed framework can generate partial shapes from a larger and more diverse pool of shapes, allowing for pairing compatible shapes from different categories, potentially originating from different datasets. This creates a more challenging scenario encompassing simultaneous non-isometric deformations and partiality, representing an important step towards more realistic settings. Furthermore, our framework introduces more variety by using different camera poses for ray casting and re-meshing shapes to different resolutions.
@@ 3.6 Semantic Missing Parts
By combining multiple categories and datasets, we encounter cases where dense correspondences cannot be reliably established between certain parts of shapes without introducing inaccuracies (e.g., matching a deer’s horn to a horse’s head). While such ambiguities can be managed when considering semantic correspondences [AEOW23], they remain problematic for dense correspondences. To address this, for the first time we create large cross-dataset correspondences with semantic missing parts annotations, as shown in Fig. 3. To this end, we modify our pipeline described in Sec. 3.1 by manually removing the semantic missing parts from the shapes and match the rest of the shapes to get the dense correspondences. We show an example in Fig. 10 in supp. mat. This creates a novel objective for shape matching methods that warrants future research.